A to bridge charges $1,00 for passenger cars and $2.75 for other vehicles. Suppose that during daytime hours, 60% of all vehicles are passenger cars. If is vehicles cross the bndge during [Hint: Let X-the number of passenger cars; then the toll revenue n(X) is a linear function of X.] ? 215 partic lar daytime panod what isthe es ting expected toll revenue?

Please shwo all work TIA
Answer #1

Out of 15 vehicles, the expected number of passenger cars are = 60% of 15 = (3/5) * 15 = 9.
Hence, the expected number of other cars are = 15 - 9 = 6.

Expected toll revenue from passenger cars = 9 * $1 = $9.
Expected toll revenue from other cars = 6 * $2.75 = $16.5.

Resulting expected toll revenue = $(9 + 16.5) = $25.5.